OpenStudy (anonymous):

In the below system, solve for y in the first equation. x + 3y = 6 2x – y = 10 Answer negative one thirdx + 2 –x + 2 negative one thirdx + 6 one thirdx + 2

OpenStudy (hoblos):

x + 3y = 6 3y = -x+6 y = -1/3 x +2

OpenStudy (hoblos):

negative one thirdx + 2